A LETTER HAS COME FROM EITHER CALCUTTA OR TATANAGAR.ONLY THE LETTER 'TA' IS VISIBLE.FIND THE PROBABILITY THAT THE LETTER HAS COME FROM TATANAGAR?

Grade:12

2 Answers

Rahul askIITiansExpert.IITR
13 years ago
if the letters TA are jointly spaced, that is to say they are in order and they are spaced near to each other then the probablity is 2/3.

Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

DO APPROVE IF THE ANSWER HELPS YOU.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Regards,

Askiitians Experts

Rahul- IIT Roorkee
Shantanu sharma
7 years ago
Solving it by Bayes theorem let P(word 1)=1/2 ,P(word 2)=1/2P(ta/word1)=2/8 & P(TA/word2)=1/7 so ...Using P(word1/ta)=2/8*1/2 / 2/8*1/2+1/7*1/2=ans solve it